OpenStudy (anonymous):

lim x=0 xsinx/2-2cosx

OpenStudy (zzr0ck3r):

can you use LHopital's?

OpenStudy (cggurumanjunath):

type the equation properly.

OpenStudy (anonymous):

lim x=0 (xsinx)/(2+2cosx)

OpenStudy (dls):

2+2cosx? are u kidding me?

OpenStudy (dls):

u wrote 2-2cosx in the question

OpenStudy (anonymous):

o i mean 2-2cosx sorry

OpenStudy (dls):

\[\Huge \frac{x sinx}{2(1+cosx)}=\frac{2xsin \frac{x}{2}\cos{\frac{x}{2}}}{2(2\cos^2 \frac{x}{2})}\] \[\Huge \frac{\cancel{2}x \{\sin \frac{x}{2}\cancel{cos{\frac{x}{2}}}}{\cancel{2}(2\cancel{\cos^2 \frac{x}{2}})}\] \[\Huge \frac{1}{2} \tan \frac{x}{2}\] =0 if x->0 Well if it is 2+2cosx

OpenStudy (dls):

okay nevermind

OpenStudy (anonymous):

yr its 2-2cosx

OpenStudy (dls):

\[\LARGE \frac{xsinx}{2-2cosx}=\frac{x(x-\frac{x^3}{6})}{2-2(1-\frac{x^2}{2})}=\frac{x^2}{x^2}=1\] Using series expansions

OpenStudy (cggurumanjunath):

final answer is 1 if you use LHopital's rule.

OpenStudy (dls):

neglect x^3/6 and other higher order terms

OpenStudy (dls):

when x->0 they tend to 0

OpenStudy (cggurumanjunath):

@dls series answer what you have given above is right answer is 1 but previously xtan(x/2) answer is zero which is wrong.

OpenStudy (dls):

read properly,he wrote 2+2cosx in the previous reply,so I changed my solution to 2+2cosx in denominator

OpenStudy (cggurumanjunath):

what ever it is ,series solution is right.

OpenStudy (ybarrap):